# 3DEnhancer: Consistent Multi-View Diffusion for 3D Enhancement

### v1

@@ 1 Introduction
The advancements in generative models [24, 19] and differentiable rendering [43] have paved the way for a new research field known as neural rendering [60]. In addition to pushing the boundaries of view synthesis [32], the generation and editing of 3D models [81, 54, 30, 35, 84, 39, 36, 34, 13] has become achievable. These methods are trained on the large-scale 3D datasets, e.g., Objaverse dataset [14], enabling fast and diverse 3D synthesis.
@@
The most relevant works to our study are 3D enhancement approaches using video diffusion models [52, 74]. While video super-resolution (VSR) models [76] can also be adapted for 3D enhancement, several challenges that make them less suitable for use as generic 3D enhancers. First, these methods are limited to enhancing 3D model reconstructions through per-instance optimization, whereas our approach can seamlessly enhance 3D outputs by integrating multi-view enhancement into the existing two-stage 3D generation frameworks (e.g., from “MVDream [54] $\to$ LGM [58]” to “MVDream $\to$ 3DEnhancer $\to$ LGM”). Second, video models often struggle with long-term consistency and fail to correct generation artifacts in 3D objects under significant viewpoint variations. Besides, video diffusion models based on temporal attention [3] face limitations in handling long videos due to memory and speed constraints. In contrast, our multi-view enhancer models texture correspondences across various views both implicitly and explicitly, by utilizing multi-view raw attention and near-view epipolar aggregation, leading to superior view consistency and higher efficiency.
In summary, we present a novel 3DEnhancer for generic 3D enhancement using multi-view denoising diffusion. Our contributions include a robust data augmentation pipeline, and the hybrid view-consistent blocks that integrate multi-view row attention and a near-view epipolar aggregation module to promote view consistency. Compared to existing enhancement methods, our multi-view 3D enhancement framework is more versatile and supports texture refinement. We conduct extensive experiments on both multi-view enhancement and per-instance optimization tasks to evaluate the model’s components. Our proposed pipeline significantly improves the quality of coarse 3D objects and consistently surpasses existing alternatives.
@@ 2 Related Work
3D Generation with Multi-view Diffusion. The success of 2D diffusion models [57, 24] has inspired their application to 3D generation. Score distillation sampling [46, 70] distills 3D from a 2D diffusion model but faces challenges like expensive optimization, mode collapse, and the Janus problem. More recent methods propose learning the 3D via a two-stage pipeline: multi-view images generation [54, 42, 53, 71] and feed-forward 3D reconstruction [27, 75, 58]. Though yielding promising results, their performance is bounded by the quality of the multi-view generative models, including the violation of strict view consistency [39] and failing to scale up to higher resolution [53]. Recent work has focused on developing more 3D-aware attention operations, such as epipolar attention [61, 28] and row-wise attention [36]. However, we find that enforcing strict view consistency remains challenging when relying solely on attention-based operations.
@@
Image and Video Super-Resolution. Image and video SR aims to improve visual quality by upscaling low-resolution content to high resolution. Research in this field has evolved from focusing on pre-defined single degradations [86, 66, 88, 37, 9, 12, 67, 5, 6, 38] (e.g., bicubic downsampling) to addressing unknown and complex degradations [82, 69, 7, 90] in real-world scenarios. To tackle real-world enhancement, some studies [82, 69, 7, 90] introduce effective degradation pipelines that simulate diverse degradations for data augmentation during training, significantly boosting performance in handling real-world cases. To achieve photorealistic enhancement, recent work has integrated various generative priors to produce detailed textures, including StyleGAN [4, 68, 79], codebook [89, 8], and the latest diffusion models [64, 91]. For instance, StableSR [64] leverages the pretrained image diffusion model, i.e., Stable Diffusion (SD) [48], for image enhancement, while Upscale-A-Video [91] further extends the diffusion model for video upscaling. Video SR networks commonly employ recurrent frame fusion [87, 67], optical flow-guided propagation [5, 6, 7, 38] or temporal attention [91] to enhance temporal consistency across adjacent frames. However, due to large spatial misalignments from viewpoint changes, these methods face challenges in establishing long-range correspondences across multi-view images, making them unsuitable for multi-view fusion for 3D. In this study, we focus on exploiting a text-to-image diffusion model to achieve robust 3D enhancement while preserving view consistency.
@@
3D Texture Enhancement. With the rapid advancement of 3D generative models [84, 2, 35, 34, 13], attention is paid to further improve 3D generation quality through a cascade 3D enhancement module. Meta 3D Gen [2, 1] proposes a UV space enhancement model to achieve sharper textures. However, training the UV-specific enhancement model requires spatially continuous UV maps, which are limited in both quantities [14] and qualities [29]. Intex [59] and SyncMVD [40] also employ UV space for generating and enhancing 3D textures. However, these techniques are specifically designed for 3D mesh with UV coordinates, making them unsuitable for other 3D representations like 3DGS [32]. Unique3D [72] and CLAY [84] apply 2D enhancement module RealESRGAN [69] directly to the generated multi-view outputs. Though straightforward, this approach risks compromising 3D consistency across the multi-view results. MagicBoost [78] introduces a 3D refinement pipeline but relies on computationally expensive SDS optimization. SuperGaussian [52] and 3DGS-Enhancer [74] propose to enhance 3D through 2D video generative priors [3, 76]. These pre-trained video models struggle to maintain long-range consistency under large viewpoint variations, making them less effective at fixing texture errors in multi-view generation.
@@ 3 Methodology
Figure 2: An overview of 3DEnhancer. By harnessing the power of generative priors, 3DEnhancer adapts a text-to-image diffusion model to a multi-view framework aimed at 3D enhancement. It is compatible with multi-view images generated by models like MVDream [54] or those rendered from coarse 3D representations, such as NeRFs [43] and 3DGS [32]. Given low-quality multi-view images along with their corresponding camera poses, 3DEnhancer aggregates multi-view information within a DiT [44] framework using row attention and epipolar aggregation modules, improving visual quality while preserving high consistency across views. Furthermore, the model supports texture-level editing via text prompts and adjustable noise levels, allowing users to correct texture errors and control the enhancement strength.
@@ 3.1 Pose-aware Encoder
Given the posed multi-view images $\mathcal{X}$, we add controllable noise to the images as an augmentation to enable controllable refinement, as described later in Sec. 3.3. To further inject camera condition for each view $v$, we follow the prior work [55, 77, 35, 58], and concatenate Plücker coordinates ${\mathbf{r}}_{\mathbf{v}}^{i}=({\mathbf{d}}^{i},{\mathbf{o}}^{i}\times{\mathbf{d}}^{i})\in\mathbb{R}^{6}$ with image RGB values ${\mathbf{x}}_{{\mathbf{v}}}^{i}\in\mathbb{R}^{3}$ along the channel dimension. Here, ${\mathbf{o}}^{i}$ and ${\mathbf{d}}^{i}$ are the ray origin and ray direction for pixel $i$ from view ${\mathbf{v}}$, and $\times$ denotes the cross product. We then send the concatenated results to a trainable pose-aware multi-view encoder $\mathcal{E}_{\bm{\psi}}$, whose outputs are injected into the pre-trained DiT through a learnable copy [83]. The visual illustration of our design is shown in Fig. 2.
@@ 3.2 View-Consistent DiT Block
The main challenge of 3D enhancement is achieving precise view consistency across generated 2D multi-view images. Multi-view diffusion methods commonly rely on multi-view attention layers to exchange information across different views, aiming to generate multiview-consistent images. A prevalent approach is extending self-attention to all views, known as dense multi-view attention [54, 42]. While effective, this method significantly raises both computational demands and memory requirements.
To further enhance the effectiveness and efficiency of inter-view aggregation, we introduce two efficient modules in the DiT blocks: multi-view row attention and near-view epipolar aggregation, as shown in Fig. 2.
Multi-view Row Attention. To enhance the noisy input views to higher resolution, e.g., $512\times 512$, efficient multi-view attention is required to facilitate cross-view information fusion. Considering the epipolar constraints [21], the 3D correspondences across multiple views always lie on the epipolar line [61, 28]. Since our diffusion denoising is performed on 16×16$\times$16 × downsampled features [10], and typical multi-view settings often involve elevation angles around $0^{\circ}$, we assume that horizontal rows approximate the epipolar line. Therefore, we adopt the special epipolar attention, specifically the multi-view row attention [36], which enables efficient information exchange among multi-view features.
@@
Near-view Epipolar Aggregation. Though multi-view attention can effectively facilitate view consistency, we observe that the attention-only operation still struggles with accurate correspondences across views. To address this issue, we incorporate explicit feature aggregation among neighboring views to ensure multi-view consistency. Specifically, given the output features $\{{\mathbf{f}}_{\mathbf{v}}\}_{{\mathbf{v}}=1}^{N}$ from the multi-view row attention layers for each posed multi-view input, we propagate features by finding near-view correspondences with epipolar line constraints. Formally, for the feature map ${\mathbf{f}}_{{\mathbf{v}}}$ corresponding to the posed image ${\mathbf{x}}_{{\mathbf{v}}}$, we calculate its correspondence map $M_{{\mathbf{v}},{\mathbf{k}}}$ with a near-view ${\mathbf{x}}_{{\mathbf{k}}}$ as follows:
@@ 3.3 Multi-view Data Augmentation
Texture Distortion. To emulate the low-quality textures and local inconsistencies found in synthesized multi-view images, we employ a texture degradation pipeline commonly used in 2D enhancement [69, 91]. This pipeline randomly applies downsampling, blurring, noise, and JPEG compression to degrade the image quality. To simulate inconsistencies between views, we adopt different stochastic degradation parameters for each view.
@@
Noise-level Control. To control the enhancement strength, we apply noise augmentation by adding controllable noise to the input multi-view images. This noise augmentation process is similar to the diffusion process in diffusion models. This approach can further enhance the model’s robustness in handling unseen artifacts [91]. Generally, higher noise levels lead to stronger refinement and regeneration. This design allows users to adjust the noise levels, balancing the enhancement and generation, as shown in Fig. 1(b).
Figure 3: Qualitative comparisons of enhancing multi-view synthesis on the Objaverse synthetic dataset. As can be seen, only 3DEnhancer can correct flowed and missing textures with view consistency. (Zoom-in for best view)
@@ 4.1 Datasets and Implementation
Figure 4: Qualitative comparisons of enhancing multi-view synthesis with RealBasicVSR[7] and Upscale-A-Video[91] on the in-the-wild dataset. Visually inspecting, 3DEnhancer yields sharp and consistent textures with intact semantics, such as the eyes of the girl in the mid row. (Zoom-in for best view)
Datasets. For training, we use the Objaverse dataset [14], specifically leveraging the G-buffer Objaverse [47], which provides diverse and high-quality renderings on Objaverse instances. We construct LQ-HQ view pairs following the augmentation pipeline outlined in Sec. 3.3 and then split the dataset into separate training and testing sets. Overall, approximately $400$K objects are used for training.
For evaluation, we use a test set containing $500$ objects from different categories within our synthesized Objaverse datasets. For each object, we render four orthogonal views from elevation angles ranging from $-5^{\circ}$ to $30^{\circ}$. Besides, we evaluate our model on the zero-shot in-the-wild dataset by selecting images from the GSO dataset [16], the output of generative image diffusion models [48], and web-sourced content. These images are then processed using several novel view synthesis methods [39, 54, 41, 36] to create our in-the-wild testing set, containing a total of 400 instances.
Implementation Details. We employ PixArt-$\Sigma$ [10], an efficient DiT model, as our backbone. Our model is trained on images with a resolution of 512 x 512. The AdamW optimizer [33] is used with a fixed learning rate of 2e-5. Our training is conducted over 10 days using 8 Nvidia A100-80G GPUs, with a batch size 256. For inference, we employ a DDIM sampler [56] with 20 steps and set the Classifier-Free Guidance (CFG) [23] scale to 4.5. During training, we randomly sample input views with azimuth angles ranging from $0^{\circ}$ to $360^{\circ}$ and elevation angles between $-5^{\circ}$ and $30^{\circ}$.
Baselines. To assess the effectiveness of our approach, we adopt two image enhancement models, RealESRGAN [69] and StableSR [64], along with two video enhancement models, RealBasicVSR [7] and Upscale-a-Video [91] as our baselines. For a fair comparison, we further fine-tune all these methods on the Objaverse dataset to minimize potential data domain discrepancies. During inference, since Real-ESRGAN, RealBasicVSR, and Upscale-a-Video by default produce images upscaled by a factor of ×4, we resize their outputs to the a uniform resolution of 512 × 512 for comparison.
Metrics. We evaluate the effectiveness of our methods on two tasks: multi-view synthesis enhancement and 3D reconstruction improvement. We employ standard metrics including PSNR, SSIM, and LPIPS [85] on our synthetic dataset, along with non-reference metrics including FID [51], Inception Score [49], and MUSIQ [31] on the in-the-wild dataset. For FID computation, we use the rendered images from Objaverse to represent the real distribution. Besides the quantitative metrics, we present comprehensive qualitative results in the paper.

### v2

@@ 1 Introduction
The advancements in generative models [24, 19] and differentiable rendering [45] have paved the way for a new research field known as neural rendering [62]. In addition to pushing the boundaries of view synthesis [33], the generation and editing of 3D models [84, 56, 31, 36, 87, 40, 37, 35, 13, 25] has become achievable. These methods are trained on the large-scale 3D datasets, e.g., Objaverse dataset [14], enabling fast and diverse 3D synthesis.
@@
The most relevant works to our study are 3D enhancement approaches using video diffusion models [54, 77]. While video super-resolution (SR) models [79] can also be adapted for 3D enhancement, several challenges that make them less suitable for use as generic 3D enhancers. First, these methods are limited to enhancing 3D model reconstructions through per-instance optimization, whereas our approach can seamlessly enhance 3D outputs by integrating multi-view enhancement into the existing two-stage 3D generation frameworks (e.g., from “MVDream [56] $\to$ LGM [60]” to “MVDream $\to$ 3DEnhancer $\to$ LGM”). Second, video models often struggle with long-term consistency and fail to correct generation artifacts in 3D objects under significant viewpoint variations. Besides, video diffusion models based on temporal attention [3] face limitations in handling long videos due to memory and speed constraints. In contrast, our multi-view enhancer models texture correspondences across various views both implicitly and explicitly, by utilizing multi-view row attention and near-view epipolar aggregation, leading to superior view consistency and higher efficiency.
In summary, we present a novel 3DEnhancer for generic 3D enhancement using multi-view denoising diffusion. Our contributions include a robust data augmentation pipeline, and the hybrid view-consistent blocks that integrate multi-view row attention and near-view epipolar aggregation modules to promote view consistency. Compared to existing enhancement methods, our multi-view 3D enhancement framework is more versatile and supports texture refinement. We conduct extensive experiments on both multi-view enhancement and per-instance optimization tasks to evaluate the model’s components. Our proposed pipeline significantly improves the quality of coarse 3D objects and consistently surpasses existing alternatives.
@@ 2 Related Work
3D Generation with Multi-view Diffusion. The success of 2D diffusion models  [59, 24] has inspired their application to 3D generation. Score distillation sampling (SDS) [48, 72] distills 3D from a 2D diffusion model but faces challenges like expensive optimization, mode collapse, and the Janus problem. More recent methods propose learning the 3D via a two-stage pipeline: multi-view images generation [56, 44, 55, 73] and feed-forward 3D reconstruction [28, 78, 60]. Though yielding promising results, their performance is bounded by the quality of the multi-view generative models, including the violation of strict view consistency [40] and failing to scale up to higher resolution [55]. Recent work has focused on developing more 3D-aware attention operations, such as epipolar attention [63, 29] and row-wise attention [37]. However, we find that enforcing strict view consistency remains challenging when relying solely on attention-based operations.
@@
Image and Video Super-Resolution. Image and video SR aim to improve visual quality by upscaling low-resolution content to high resolution. Research in this field has evolved from focusing on pre-defined single degradations [89, 68, 92, 38, 9, 12, 69, 5, 6, 39] (e.g., bicubic downsampling) to addressing unknown and complex degradations [85, 71, 7] in real-world scenarios. To tackle real-world enhancement, some studies [85, 71, 7, 94] introduce effective degradation pipelines that simulate diverse degradations for data augmentation during training, significantly boosting performance in handling real-world cases. To achieve photorealistic enhancement, recent work has integrated various generative priors to produce detailed textures, including StyleGAN [4, 70, 82], codebook [93, 8], and the latest diffusion models [66, 95]. For instance, StableSR [66] leverages the pretrained image diffusion model, i.e., Stable Diffusion (SD) [50], for image enhancement, while Upscale-A-Video [95] further extends the diffusion model for video upscaling. Video SR networks commonly employ recurrent frame fusion [91, 69], optical flow-guided propagation [5, 6, 7, 39] or temporal attention [95] to enhance temporal consistency across adjacent frames. However, due to large spatial misalignments from viewpoint changes, these methods face challenges in establishing long-range correspondences across multi-view images, making them unsuitable for multi-view fusion for 3D. In this study, we focus on exploiting a image diffusion model to achieve robust 3D enhancement while preserving view consistency.
@@
3D Texture Enhancement. With the rapid advancement of 3D generative models [87, 2, 36, 35, 13], attention is paid to further improve 3D generation quality through a cascade 3D enhancement module. Meta 3D Gen [2, 1] proposes a UV space enhancement model to achieve sharper textures. However, training the UV-specific enhancement model requires spatially continuous UV maps, which are limited in both quantities [14] and qualities [30]. Intex [61] and SyncMVD [42] also employ UV space for generating and enhancing 3D textures. However, these techniques are specifically designed for 3D mesh with UV coordinates, making them unsuitable for other 3D representations like 3DGS [33]. Unique3D [74] and CLAY [87] apply 2D enhancement module RealESRGAN [71] directly to the generated multi-view outputs. Though straightforward, this approach risks compromising 3D consistency across the multi-view results. MagicBoost [81] introduces a 3D refinement pipeline but relies on computationally expensive SDS optimization. Deceptive-NeRF/3DGS [41] uses an image diffusion model to generate high-quality pseudo-observations for novel views but requires a few accurately captured sparse views as key inputs. SuperGaussian [54] and 3DGS-Enhancer [77] propose to enhance 3D through 2D video generative priors [3, 79]. These pre-trained video models struggle to maintain long-range consistency under large viewpoint variations, making them less effective at fixing texture errors in multi-view generation.
@@ 3 Methodology
Figure 2: An overview of 3DEnhancer. By harnessing generative priors, 3DEnhancer adapts a text-to-image diffusion model to a multi-view framework aimed at 3D enhancement. It is compatible with multi-view images generated by models like MVDream [56] or those rendered from coarse 3D representations like NeRFs [45] and 3DGS [33]. Given LQ multi-view images along with their corresponding camera poses, 3DEnhancer aggregates multi-view information within a DiT [46] framework using row attention and epipolar aggregation modules, improving visual quality while preserving consistency across views. Furthermore, the model supports texture-level editing via text prompts and adjustable noise levels, allowing users to correct texture errors and control the enhancement strength.
@@ 3.1 Pose-aware Encoder
Given the posed multi-view images $\mathcal{X}$, we add controllable noise to the images as an augmentation to enable controllable refinement, as described later in Sec. 3.3. To further inject camera condition for each view $v$, we follow the prior work [57, 80, 36, 60], and concatenate Plücker coordinates ${\mathbf{r}}_{\mathbf{v}}^{i}=({\mathbf{d}}^{i},{\mathbf{o}}^{i}\times{\mathbf{d}}^{i})\in\mathbb{R}^{6}$ with image RGB values ${\mathbf{x}}_{{\mathbf{v}}}^{i}\in\mathbb{R}^{3}$ along the channel dimension. Here, ${\mathbf{o}}^{i}$ and ${\mathbf{d}}^{i}$ are the ray origin and ray direction for pixel $i$ from view ${\mathbf{v}}$, and $\times$ denotes the cross product. We then send the concatenated results to a trainable pose-aware multi-view encoder $\mathcal{E}_{\bm{\psi}}$, whose outputs are injected into the pre-trained DiT through a learnable copy [86].
@@ 3.2 View-Consistent DiT Block
The main challenge of 3D enhancement is achieving precise view consistency across generated 2D multi-view images. Multi-view diffusion methods commonly rely on multi-view attention layers to exchange information across different views, aiming to generate multiview-consistent images. A prevalent approach is extending self-attention to all views, known as dense multi-view attention [56, 44]. While effective, this method significantly raises both computational demands and memory requirements. To further enhance the effectiveness and efficiency of inter-view aggregation, we introduce two efficient modules in the DiT blocks: multi-view row attention and near-view epipolar aggregation, as shown in Fig. 2.
Multi-view Row Attention. To enhance the noisy input views to higher resolution, e.g., $512\times 512$, efficient multi-view attention is required to facilitate cross-view information fusion. Considering the epipolar constraints [21], the 3D correspondences across views always lie on the epipolar line [63, 29]. Since our diffusion denoising is performed on 16×16$\times$16 × downsampled features [10], and typical multi-view settings often involve elevation angles around $0^{\circ}$, we assume that horizontal rows approximate the epipolar line. Therefore, we adopt the special epipolar attention, specifically the multi-view row attention [37], enabling efficient information exchange among multi-view features.
@@
Near-view Epipolar Aggregation. Though multi-view attention can effectively facilitate view consistency, we observe that the attention-only operation still struggles with accurate correspondences across views. To address this issue, we incorporate explicit feature aggregation among neighboring views to ensure multi-view consistency. Specifically, given the output features $\{{\mathbf{f}}_{\mathbf{v}}\}_{{\mathbf{v}}=1}^{N}$ from the multi-view row attention layers for each posed multi-view input, we propagate features by finding near-view correspondences with epipolar line constraints. Formally, for the feature map ${\mathbf{f}}_{{\mathbf{v}}}$ corresponding to the posed image ${\mathbf{x}}_{{\mathbf{v}}}$, we calculate its correspondence map $M_{{\mathbf{v}},{\mathbf{k}}}$ with the near views ${\mathbf{k}}$ as follows:
@@ 3.3 Multi-view Data Augmentation
Figure 3: Qualitative comparisons of enhancing multi-view synthesis on the Objaverse synthetic dataset. As can be seen, only 3DEnhancer can correct flowed and missing textures with view consistency.
Texture Distortion. To emulate the low-quality textures and local inconsistencies found in synthesized multi-view images, we employ a texture degradation pipeline commonly used in 2D enhancement [71, 95]. This pipeline randomly applies downsampling, blurring, noise, and JPEG compression to degrade the image quality.
@@
Noise-level Control. To control the enhancement strength, we apply noise augmentation by adding controllable noise to the input multi-view images. This noise augmentation process is similar to the diffusion process in diffusion models. This approach can further enhance the model’s robustness in handling unseen artifacts [95].
@@ 4.1 Datasets and Implementation
Datasets. For training, we use the Objaverse dataset [14], specifically leveraging the G-buffer Objaverse [49], which provides diverse renderings on Objaverse instances. We construct LQ-HQ view pairs following the augmentation pipeline outlined in Sec. 3.3 and then split the dataset into separate training and test sets. Overall, approximately $400$K objects are used for training. For each object, we randomly sample 4 input views with azimuth angles ranging from $0^{\circ}$ to $360^{\circ}$ and elevation angles between $-5^{\circ}$ and $30^{\circ}$.
For evaluation, we use a test set containing $500$ objects from different categories within our synthesized Objaverse datasets. We further evaluate our model on the zero-shot in-the-wild dataset by selecting images from the GSO dataset [16], image diffusion model outputs [50], and web-sourced content. These images are then processed using several novel view synthesis methods [40, 56, 43, 37] to create our in-the-wild test set, containing a total of 400 instances.
Implementation Details. We employ PixArt-$\Sigma$ [10], an efficient DiT model, as our backbone. Our model is trained on images with a resolution of 512 x 512. The AdamW optimizer [34] is used with a fixed learning rate of 2e-5. Our training is conducted over 10 days using 8 Nvidia A100-80G GPUs, with a batch size 256. For inference, we employ a DDIM sampler [58] with 20 steps and set the Classifier-Free Guidance (CFG) [23] scale to 4.5.
Figure 4: Qualitative comparisons of enhancing multi-view synthesis with RealBasicVSR[7] and Upscale-A-Video[95] on the in-the-wild dataset. Visually inspecting, 3DEnhancer yields sharp and consistent textures with intact semantics, such as the eyes of the girl.
Baselines. To assess the effectiveness of our approach, we adopt two image enhancement models, RealESRGAN [71] and StableSR [66], along with two video enhancement models, RealBasicVSR [7] and Upscale-a-Video [95] as our baselines. For a fair comparison, we further fine-tune all these methods on the Objaverse dataset to minimize potential data domain discrepancies. During inference, since Real-ESRGAN, RealBasicVSR, and Upscale-a-Video by default produce images upscaled by a factor of ×4, we resize their outputs to a uniform resolution of 512 × 512 for comparison.
Metrics. We evaluate the effectiveness of our methods on two tasks: multi-view synthesis enhancement and 3D reconstruction improvement. We employ standard metrics including PSNR, SSIM, and LPIPS [88] on our synthetic dataset, along with non-reference metrics including FID [53], Inception Score [51], and MUSIQ [32] on the in-the-wild dataset. For FID computation, we use the rendered images from Objaverse to represent the real distribution.